# Gamma-Ray Observations of Galaxy Clusters Strongly Constrain Dark Matter Annihilation in Prompt Cusps

### v1

@@ III Data Analysis
The Large Area Telescope (LAT), onboard the Fermi $\gamma$-ray observatory, detects $\gamma$ rays with energies spanning 20 MeV– 500 GeV. Its field of view (FoV) covers $\sim$20of the sky at all times, allowing full-sky coverage at a cadence of about 3 hours. Its source localization accuracy reaches a few arcminutes. The instrument paper provides a comprehensive overview of the LAT instrument and its performance Atwood et al. (2009).
We analyze 15 years of Pass 8 reprocessed data covering the period from August 4, 2008, to November 1, 2023, obtained from the Fermi Science Support Center (FSSC) website [^1] https://fermi.gsfc.nasa.gov/ssc/data/, accessed on January 19, 2025.. All data are analyzed using the fermipy software package (v. 1.2.0) Wood et al. (2017), which relies on FermiTools (v. 2.2.0) [^2] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/, accessed on January 19, 2025. for likelihood-based modeling and fitting. We select photons from the P8R3_SOURCE_V3 event class with energies between 500 MeV and 500 GeV, spaced in 8 logarithmic bins in each decade in energy. We apply the standard models for the Galactic diffuse emission (gll_iem_v07) and the isotropic diffuse background (iso_P8R3_SOURCE_V3), as provided by the Fermi-LAT Collaboration. For each source, we define a 20$\times$20 region of interest (RoI), binned at 0.2 and centered on the cluster’s dynamical center. We model both the point and extended sources within the RoI using the 4FGL-DR3 catalog (gll_psc_v29), extending to 25$\times$25 to account for photon leakage from the LAT point spread function. We introduce a zenith angle cut of 90 to account for possible contamination from the Earth’s limb and apply the standard good time interval (GTI) selection cuts.
@@
After constructing the model, we perform a binned likelihood analysis for each source. The likelihood function compares models with and without an associated $\gamma$-ray source, quantifying the goodness of fit using the Test Statistic (TS), defined as $TS=-2~{}{\rm log}\left(\mathcal{L}_{0}/\mathcal{L}_{1}\right)$, where $\mathcal{L}_{0}$ represents the likelihood of the null hypothesis (no signal), and $\mathcal{L}_{1}$ is the likelihood of the alternative hypothesis (signal present). We perform an extended-source analysis using the dark-matter signal templates described in Sec. II for the seven clusters from Table 1. We model the dark matter profile from each cluster out to 6 from the cluster center, in order to model the entire dark matter contribution (see Figure 2). For each cluster, we use the gta.sed() module in fermipy to estimate the source flux by fitting the spectrum in each energy bin assuming a power-law shape with an index of 2. This method also produces a spectral energy distribution (SED) curve, and energy flux $E^{2}\mathrm{d}N/\mathrm{d}E$, along with a corresponding delta log-likelihood profile, $\Delta\mathcal{L}\left(\mathrm{d}\Phi_{\gamma}/\mathrm{d}E,E\right)$. Once the flux (or upper limit) is determined, we translate the SEDs into dark matter constraints. The $\gamma$-ray flux from dark matter annihilation is:
@@
where $\langle\sigma v\rangle$ is the velocity-weighted annihilation cross-section, $m_{\text{DM}}$ is the mass of the dark matter particle, and $\mathrm{d}N/\mathrm{d}E$ is the spectrum of $\gamma$ rays produced per annihilation event (depending on the dark matter mass and annihilation channel) Cirelli et al. (2011). Finally, we are then able to constrain the dark matter parameter space by calculating $TS$ values for each combination of cross-section and mass,
$TS\left(\langle\sigma v\rangle,m_{\text{DM}}\right)=-2~{}{\rm log}\frac{\mathcal{L}_{0}}{\mathcal{L}\left(\langle\sigma v\rangle,m_{\text{DM}}\right)}$ (15)
where $\mathcal{L}_{0}$ is the likelihood of the null hypothesis (no dark matter) and $\mathcal{L}\left(\langle\sigma v\rangle,M_{\chi}\right)$ represents the likelihood of the model with dark matter. We sample 1,000 logarithmically spaced dark matter masses between 10 GeV–10 TeV and 60,000 $\langle\sigma v\rangle$ values logarithmically spaced within 5 orders of magnitude around the thermal relic cross-section ($3\times 10^{-26}$ cm3/s) in both directions. We use J factors from Fig. 3 and annihilation channels that produce $b\bar{b}$, $\tau^{+}\tau^{-}$, and $W^{+}W^{-}$ final states. By stepping through the values of $\langle\sigma v\rangle$, we determine the cross-section value corresponding to a 95% confidence level (CL) limit, defined as the point where the $TS$ value decreases by 2.71.
@@ IV Results
We find no evidence for a spatially extended $\gamma$-ray signal in any of the seven galaxy clusters listed in Table 1. Thus, we place upper limits on the dark matter annihilation cross section using 15 years of Fermi-LAT data.
The Virgo cluster, unsurprisingly, provides the strongest constraints among our targets, primarily due to its large annihilation $J$-factor (Fig. 3). Figure 1 compares our Virgo limits for the $b\bar{b}$ annihilation channel to other relevant benchmarks, showing that we rule out dark matter masses below $\sim$200 GeV at a 95% CL. Below $\sim$1 TeV, our Virgo limit is stronger by about an order of magnitude compared to the stacking of dwarf spheroidal galaxies, while at higher masses (up to $\sim$10 TeV) it remains competitive McDaniel et al. (2024). Finally, our annihilation limits are about an order of magnitude more stringent than previous limits from galaxy clusters Di Mauro et al. (2023) and the broader large-scale structure Bartlett et al. (2022), which relied on conventional subhalo models that did not account for prompt cusps.
@@ Acknowledgements
MC and TL acknowledge support from the Swedish Research Council under contract 2022-04283 and the Swedish National Space Agency under contract 2023-00242.
@@ Appendix A Pericenter distribution
×$\hphantom{=}\times\delta_{\mathrm{D}}\!\left\{\Phi(r)-\Phi(r_{\mathrm{p}})-\left(\sin^{2}\theta\frac{r^{2}}{r_{\mathrm{p}}^{2}}-1\right)\frac{v^{2}}{2}\right\}$
×$\hphantom{=}\times\left|\sin^{2}\theta\frac{r^{2}}{r_{\mathrm{p}}^{3}}v^{2}-\Phi^{\prime}(r_{\mathrm{p}})\right|\mathrm{d}r_{\mathrm{p}},$ (18)
@@
×$\hphantom{=}\times\theta_{\mathrm{H}}\!\left[v^{2}-2r_{\mathrm{p}}^{2}\frac{\Phi(r)-\Phi(r_{\mathrm{p}})}{r^{2}-r_{\mathrm{p}}^{2}}\right]$
×$\hphantom{=}\times\frac{r_{\mathrm{p}}}{rv}\frac{|v^{2}+2(\Phi(r)-\Phi(r_{\mathrm{p}}))-r_{\mathrm{p}}\Phi^{\prime}(r_{\mathrm{p}})|}{\sqrt{(r^{2}-r_{\mathrm{p}}^{2})v^{2}-2r_{\mathrm{p}}^{2}(\Phi(r)-\Phi(r_{\mathrm{p}}))}}\mathrm{d}r_{\mathrm{p}},$ (19)
@@
×$\hphantom{=}\times\frac{|v^{2}+2(\Phi(r)-\Phi(r_{\mathrm{p}}))-r_{\mathrm{p}}\Phi^{\prime}(r_{\mathrm{p}})|}{\sqrt{(r^{2}-r_{\mathrm{p}}^{2})v^{2}-2r_{\mathrm{p}}^{2}(\Phi(r)-\Phi(r_{\mathrm{p}}))}},$ (20)

### v2

@@ III Data Analysis
The Large Area Telescope (LAT), onboard the Fermi $\gamma$-ray observatory, detects $\gamma$ rays with energies spanning 20 MeV– 500 GeV. Its field of view (FoV) covers $\sim$20% of the sky at all times, allowing full-sky coverage at a cadence of about 3 hours. Its source localization accuracy reaches a few arcminutes. The instrument paper provides a comprehensive overview of the LAT instrument and its performance Atwood et al. (2009).
We analyze 15 years of Pass 8 reprocessed data covering the period from August 4, 2008, to November 1, 2023, obtained from the Fermi Science Support Center (FSSC) website [^3] https://fermi.gsfc.nasa.gov/ssc/data/, accessed on January 19, 2025.. All data are analyzed using the fermipy software package (v. 1.2.0) Wood et al. (2017), which relies on FermiTools (v. 2.2.0) [^4] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/, accessed on January 19, 2025. for likelihood-based modeling and fitting. We select photons from the P8R3_SOURCE_V3 event class with energies between 500 MeV and 500 GeV, spaced in 8 logarithmic bins in each decade in energy. We apply the standard models for the Galactic diffuse emission (gll_iem_v07) and the isotropic diffuse background (iso_P8R3_SOURCE_V3), as provided by the Fermi-LAT Collaboration. For each source, we define a $\times20$ region of interest (RoI), binned at 0.2 and centered on the cluster’s dynamical center. We model both the point and extended sources within the RoI using the 4FGL-DR4 catalog (gll_psc_v31), extending to $\times25$ to account for photon leakage from the LAT point spread function. We introduce a zenith angle cut of 90 to account for possible contamination from the Earth’s limb and apply the standard good time interval (GTI) selection cuts.
@@
After constructing the model, we perform a binned likelihood analysis for each source. The likelihood function compares models with and without an associated $\gamma$-ray source, quantifying the goodness of fit using the Test Statistic (TS), defined as $TS=-2~{\rm log}\left(\mathcal{L}_{0}/\mathcal{L}_{1}\right)$, where $\mathcal{L}_{0}$ represents the likelihood of the null hypothesis (no signal), and $\mathcal{L}_{1}$ is the likelihood of the alternative hypothesis (signal present). We perform an extended-source analysis using the dark-matter signal templates described in Sec. II for the seven clusters from Table 1. We model the dark matter profile from each cluster out to 6 from the cluster center, in order to model the entire dark matter contribution (see Figure 2). For each cluster, we use the gta.sed() module in fermipy to estimate the source flux by fitting the spectrum in each energy bin assuming a power-law shape with an index of 2. This method also produces a spectral energy distribution (SED) curve, and energy flux $E^{2}\mathrm{d}N/\mathrm{d}E$, along with a corresponding delta log-likelihood profile, $\Delta\mathcal{L}\left(\mathrm{d}\Phi_{\gamma}/\mathrm{d}E,E\right)$.
Once the flux (or upper limit) is determined, we translate the SEDs into dark matter constraints. The $\gamma$-ray flux from dark matter annihilation is:
@@
where $\mathrm{d}N/\mathrm{d}E$ is the spectrum of $\gamma$-rays produced per annihilation event, obtained from PPPC4DM Cirelli et al. (2011) and dependent on both the dark matter mass and annihilation channel. We constrain the dark matter parameter space by calculating $TS$ values for each combination of cross-section and mass,
$TS\left(\langle\sigma v\rangle,m_{\text{DM}}\right)=-2~{\rm log}\frac{\mathcal{L}_{0}}{\mathcal{L}\left(\langle\sigma v\rangle,m_{\text{DM}}\right)}.$ (16)
We sample 1,000 logarithmically spaced dark matter masses between 10 GeV–10 TeV and 60,000 $\langle\sigma v\rangle$ values logarithmically spaced within 5 orders of magnitude around the thermal relic cross-section ($3\times 10^{-26}$ cm3/s) in both directions. We use J factors from Fig. 3 and annihilation channels that produce $b\bar{b}$, $\tau^{+}\tau^{-}$, and $W^{+}W^{-}$ final states. By stepping through the values of $\langle\sigma v\rangle$, we determine the cross-section value corresponding to a 95% confidence level (CL) limit, defined as the point where the $TS$ value decreases by 2.71. Detailed analysis results including TS maps, SEDs, and individual cluster results are presented in Appendix B.
@@ IV Results
We find no evidence for a spatially extended $\gamma$-ray signal in any of the seven galaxy clusters listed in Table 1. We note that, however, when fitting dark matter annihilation spectra to the observed flux limits, several clusters show modest statistical preferences for DM models over background-only scenarios (see Appendix B for detailed individual cluster analysis). Thus, we place upper limits on the dark matter annihilation cross section using 15 years of Fermi-LAT data.
The Virgo cluster, unsurprisingly, provides the strongest constraints among our targets, primarily due to its large annihilation $J$-factor (Fig. 3). Figure 1 compares our Virgo limits for the $b\bar{b}$ annihilation channel to other relevant benchmarks, showing that we rule out dark matter masses below $\sim$200 GeV at a 95% CL. Below $\sim$1 TeV, our Virgo limit is stronger by about an order of magnitude compared to the stacking of dwarf spheroidal galaxies, while at higher masses (up to $\sim$10 TeV) it remains competitive McDaniel et al. (2024). Finally, our annihilation limits are about an order of magnitude more stringent than previous limits from galaxy clusters Di Mauro et al. (2023); Manna and Desai (2024) and the broader large-scale structure Bartlett et al. (2022), which relied on conventional subhalo models that did not account for prompt cusps.
@@ V Conclusion
We note that galaxy clusters can produce $\gamma$-ray emission from cosmic-ray interactions with the intracluster medium, creating potential astrophysical backgrounds for dark matter searches Pinzke and Pfrommer (2010); Huber et al. (2013); Hussain et al. (2023). While a detailed treatment of these cosmic-ray backgrounds is beyond the scope of this study, our constraints remain robust because prompt cusp annihilation produces distinctive spatial signatures that differ from the extended emission expected from cosmic-ray processes.
@@ Acknowledgements
MC and TL acknowledge support from the Swedish Research Council under contract 2022-04283 and the Swedish National Space Agency under contract 2023-00242. TL also acknowledges sabbatical support from the Wenner-Gren foundation under contract SSh2024-0037.
@@ Appendix A Pericenter distribution
$\hphantom{=}\times\delta_{\mathrm{D}}\!\left\{\Phi(r)-\Phi(r_{\mathrm{p}})-\left(\sin^{2}\theta\frac{r^{2}}{r_{\mathrm{p}}^{2}}-1\right)\frac{v^{2}}{2}\right\}$
$\hphantom{=}\times\left|\sin^{2}\theta\frac{r^{2}}{r_{\mathrm{p}}^{3}}v^{2}-\Phi^{\prime}(r_{\mathrm{p}})\right|\mathrm{d}r_{\mathrm{p}},$ (19)
@@
$\hphantom{=}\times\theta_{\mathrm{H}}\!\left[v^{2}-2r_{\mathrm{p}}^{2}\frac{\Phi(r)-\Phi(r_{\mathrm{p}})}{r^{2}-r_{\mathrm{p}}^{2}}\right]$
$\hphantom{=}\times\frac{r_{\mathrm{p}}}{rv}\frac{|v^{2}+2(\Phi(r)-\Phi(r_{\mathrm{p}}))-r_{\mathrm{p}}\Phi^{\prime}(r_{\mathrm{p}})|}{\sqrt{(r^{2}-r_{\mathrm{p}}^{2})v^{2}-2r_{\mathrm{p}}^{2}(\Phi(r)-\Phi(r_{\mathrm{p}}))}}\mathrm{d}r_{\mathrm{p}},$ (20)
@@
$\hphantom{=}\times\frac{|v^{2}+2(\Phi(r)-\Phi(r_{\mathrm{p}}))-r_{\mathrm{p}}\Phi^{\prime}(r_{\mathrm{p}})|}{\sqrt{(r^{2}-r_{\mathrm{p}}^{2})v^{2}-2r_{\mathrm{p}}^{2}(\Phi(r)-\Phi(r_{\mathrm{p}}))}},$ (21)
@@
Appendix B Fermi-LAT Analysis
Here, we provide supplementary details from our fermipy-based $\gamma$-ray analysis of the seven galaxy clusters listed in Table 1. The analysis steps are summarized below, followed by individual cluster results.
[…]
  • Data Selection: We select photon events from the P8R3_SOURCE_V3 event class.
  • Quality Cuts: We apply standard quality cuts, including a zenith angle cut of 90 to mitigate contamination from the Earth’s limb.
  • Binning: We bin the data spatially with a pixel size of 0.2, and in energy with 8 logarithmically spaced bins per decade.
[…]
  • Source Modeling: All confirmed point and extended sources were included in our background model using the 4FGL-DR4 parameters Abdollahi et al. (2022). We model sources within a 25 radius of the RoI center to account for contributions from the tails of the Fermi-LAT point spread function.
  • Diffuse Emission Models: We incorporate the standard Galactic diffuse emission model (gll_iem_v07.fits) and the isotropic diffuse background model appropriate for the event class (iso_P8R3_SOURCE_V3_v1.txt).
We model each galaxy cluster as an extended source using a spatial template derived from the prompt cusp annihilation profiles (Section II). This component is added to the model as a SpatialMap, with a fixed spectral index of 2.0 and a free normalization parameter.
After building the sky model, we optimize source parameters by freeing the normalizations (norm) of all catalog sources within $5^{\circ}$ of the RoI center and with TS >>25, as well as all spectral parameters of catalog sources within $7^{\circ}$ of the RoI center and with TS >> 500. The normalization parameters of the diffuse components (Galactic and isotropic) are also left free to vary. To reduce computation time and prevent instability in the fitting procedure, we disable energy dispersion corrections for these diffuse components. This is justified because energy dispersion effects for smooth, extended backgrounds typically result in negligible improvement in fit quality. For the cluster templates and point sources within the analysis region, energy dispersion remains enabled.
We then perform a maximum likelihood fit using gta.fit(), which adjusts all the free parameters to best match the data. We extract signal or spectral upper limits using the SED likelihood procedure, gta.sed().
B.1 Dark Matter Analysis
We convert the measured $\gamma$-ray flux limits to dark matter constraints using the expression in Eq. (15) and then, for each combination of dark matter mass and velocity-weighted annihilation cross-section, we calculate the TS in Eq. (16). We consider three annihilation channels ($b\bar{b}$, $\tau^{+}\tau^{-}$, and $W^{+}W^{-}$) using $\gamma$-ray spectra from PPPC4DM Cirelli et al. (2011) with electroweak corrections enabled.
TS Map DM Constraints SED
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Comprehensive analysis results for all seven galaxy clusters. Left column: Test statistic ($\sqrt{\text{TS}}$) maps showing no significant $\gamma$-ray excesses in any cluster region. The white cross marks each cluster’s dynamical center. Middle column: Dark matter annihilation cross-section constraints as a function of mass for the $b\bar{b}$ channel, with contours showing 68%, 95%, and 99.7% confidence levels. White solid contours indicate the strongest exclusion regions. Right column: Spectral energy distributions showing flux upper limits (black arrows) and likelihood profiles (colored bands) across energy bins from 500 MeV to 500 GeV. The blue curve in Virgo, Fornax, and Coma panels shows the expected spectrum for the best-fit dark matter model.
TS Map DM Constraints SED
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7 continued:
B.2 Fermi-LAT results
Figure 7 presents comprehensive Fermi-LAT analysis results for all seven galaxy clusters. All results shown here are computed using the fixed $J$-factor values listed in Table 1 (this is in contrast to our main results shown in Fig. 5 which account for the $m_{\mathrm{DM}}$-dependent $J$-factors shown in Fig. 3). We find no statistically significant $\gamma$-ray sources at cluster locations reported in Table 1. The way to interpret Fig. 7 is as follows:
  • Left column (TS Maps): Test statistic ($\sqrt{\text{TS}}$) maps assess $\gamma$-ray emission significance across each cluster’s RoI. These maps assume a power-law spectral model with fixed photon index ($\Gamma=2.0$) but free normalization, serving as a proxy for the strength of $\gamma$-ray emission at each spatial location covered by the DM annihilation map derived in Sec. II.1. The white cross marks each cluster’s dynamical center, and the color scale indicates the square root of the TS, which approximates the detection significance in units of $\sigma$. Values above $\sqrt{\text{TS}}\gtrsim 3$ would indicate potential $\gamma$-ray sources warranting further investigation.
  • Middle column (DM Constraints): DM parameter space constraints show the TS as a function of dark matter mass ($m_{\text{DM}}$) and velocity-averaged annihilation cross-section ($\langle\sigma v\rangle$) in Eq. 16, for the $b\bar{b}$ annihilation channel, computed using each cluster’s specific $J$-factor from Table 1. Here, TS follows a $\chi^{2}$ distribution with 2 degrees of freedom (DM mass and velocity-weightet annihilation cross-section), and the contours indicate 68%, 95%, and 99.7% CL. White solid contours delineate the strongest exclusion regions for the assumed $J$-factor values.
  • Right column (SEDs): Spectral energy distributions display the differential $\gamma$-ray flux ($E^{2}\text{d}N/\text{d}E$) as a function of energy from 500 MeV to 500 GeV. Black arrows indicate 95% confidence ULs for energy bins where no significant emission is detected (TS $\lesssim 9$), while error bars show flux measurements with uncertainties for bins with modest evidence (TS $\gtrsim 9$). The pink colored likelihood bands represent $\Delta\log\mathcal{L}$ profiles across energy, where TS (middle column) $2\Delta\log\mathcal{L}$, summed over all energy bins. For clusters showing elevated TS values, the blue curves indicate the best-fit DM annihilation spectra for the assumed $J$-factor and the $b\bar{b}$ annihilation channel.
B.3 Notes on individual clusters
Virgo Cluster.
The Virgo region contains multiple $\gamma$-ray components. M87 (4FGL J1230.2+1223) is an established point source with average flux $F_{\gamma}\sim 1.6470\times 10^{-9}$ ph cm-2 s-1 ($1~\text{GeV}<E<100$ GeV) and photon index $\Gamma=2.06\pm 0.04$ Abdollahi et al. (2022). It is also a flaring source: the EHT Collaboration detected the first $\gamma$-ray flare in over a decade during their 2018 EHT campaign Algaba et al. (2024). Additionally, Ackermann et al. reported a statistically significant extended emission with uniformly emitting disk profile (radius = 3), though its significance strongly depends on the adopted interstellar emission model and “is most likely an artifact of our incomplete description of the IEM in this region.”  Ackermann et al. (2015c)
[…]
NGC 4636.
NGC 4636 is a giant elliptical galaxy within the Virgo supercluster. Our analysis finds no significant $\gamma$-ray excess. A TS $\sim$ 30 preference around $m_{\text{DM}}\sim$ 441 GeV for the $b\bar{b}$ annihilation channel appears in our SED analysis. This represents the highest-significance preference for a DM annihilation spectral model in our cluster sample, though we note that the signal remains consistent with astrophysical origins (especially considering the crowded environment of NGC 4636).
M49.
M49 is a giant elliptical galaxy located within the Virgo supercluster. Our analysis reveals no significant $\gamma$-ray excess associated with its expected DM template. The maximum TS associated with the DM parameter space is $\sim$ 6. The SED analysis yields robust upper limits across all energy bins.
Fornax (NGC 1399).
The Fornax cluster, centered on the elliptical galaxy NGC 1399, shows no excess $\gamma$-ray emission above background expectation. When fitting the DM annihilation spectrum to the observed flux limits, the data exhibit a modest statistical preference for DM model, with a peak TS $\sim 12$ (corresponding to $\sim 3.5\sigma$). The SED analysis identifies a best-fit DM mass of $\sim$ 962 GeV for annihilation into the $b\bar{b}$ channel.
Hydra Cluster.
Our analysis does not detect any significant $\gamma$-ray excess, with DM spectral fitting yielding a maximum TS value of $\sim$7 . The SED analysis confirms that flux upper limits are consistent with background expectations across all energies.
Coma Cluster.
The Coma cluster has confirmed $\gamma$-ray emission from multiple independent analysis Xi et al. (2018); Adam et al. (2021); Baghmanyan et al. (2022). Adam et al. confirm detection with TS $\sim$ 27 using a binned likelihood approach, with $\gamma$-ray morphology elongated in the northeast–southwest direction coincident with a radio halo Adam et al. (2021). Additionally, a point source 4FGL J1256.9+2736 about 0.8 from the Coma’s dynamical center has a TS $\sim$ 17 in the 4FGL catalog Abdollahi et al. (2022), which could either be astrophysical contamination or correspond to the peak of diffuse cluster emission.
[…]
Centaurus Cluster.
Centaurus hosts extended $\gamma$-ray emission from giant lobes of Centaurus, with emission extending beyond radio image to energies up to 30 GeV Sun et al. (2016). We find no statistically significant excess in our analysis, with a DM model maximum TS value of $\sim 8$. Similar to other clusters in our sample, the SED analysis returns flux upper limits comparable with background fluctuations across the energy range considered.
B.4 Concluding remarks
[…]
These TS values are not necessarily consistent with a DM interpretation when considered collectively. Instead, they likely originate from known astrophysical processes Ackermann et al. (2015c); Abdollahi et al. (2022); Algaba et al. (2024); Xi et al. (2018); Adam et al. (2021); Baghmanyan et al. (2022); Sun et al. (2016): Coma’s signal can correlate with its confirmed radio halo and diffuse emission. Notably, NGC 4636, Virgo, and M49 are all members of the Virgo supercluster, suggesting that any large-scale astrophysical processes affecting one system could extend to the others and provide a common origin for observed $\gamma$-ray activity. Critically, all putative DM signals from other clusters are excluded by the robust upper limits derived from Virgo—our best characterized system with the largest $J$-factor and proximity.